What is the velocity of a beam of electrons that goes undeflected when moving perpendicular to an electric and magnetic fields. E⃗ and B⃗ are also perpendicular to each other and have magnitudes

4900 V/m and 3.9×10−3 T , respectively.
I got the answer to this part: 1.26*10^6 m/s.
I need help with the second part:
What is the radius of the electron orbit if the electric field is turned off?
Answer #1

Force of the magnetic field will be the Centripetal Force.
F = qvB
F = mv^2/r

Equating each other,
qvB = mv^2/r
r = m*v/qB
r = (9.1*10^-31 * 1;26*10^6)/(1.6*10^-19 * 3.9*10^-3)
r = 0.0018375 m
Radius of the orbit, r = 1.84 mm